OpenStudy (anonymous):

zero exponent property (2x^2y^-3/-12z^-1)^0

OpenStudy (anonymous):

Know that a^0 = 1

OpenStudy (anonymous):

so all equals 1?

OpenStudy (anonymous):

(2x^2y^-3/-12z^-1)^0 = 1

OpenStudy (anonymous):

can i show u another problem?

OpenStudy (anonymous):

yep, bearsams

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

(-9x^-3z^4/6y^6)^0 (-2x^2)^3

OpenStudy (anonymous):

Anything to the zero power is just one. It's the Law — the Laws of Exponents

OpenStudy (anonymous):

Are they two different terms?

OpenStudy (anonymous):

yah but 1 problem

OpenStudy (anonymous):

Then (-9x^-3z^4/6y^6)^0 (-2x^2)^3 = 1{(-2 x -2 x -2)(x^2)(x^2)(x^2)} = -8x^6

OpenStudy (anonymous):

Are you okay with that problem?

OpenStudy (anonymous):

im sorry but i dont know how to do that

OpenStudy (anonymous):

In indices, \[a^m \times a^n= a^\left( m+n \right)\]

OpenStudy (anonymous):

This implies \[(-2)^3 = -2\times-2\times-2=-8\]

OpenStudy (anonymous):

Also, \[x^2\times x^2\times x^2 = x^\left( 2+2+2 \right)\] which equals \[x^6\]